OpenStudy (anonymous):

A gallon of gas cost 30 cents in 1960 and 70 cents in 1977. What was the percent increase?

OpenStudy (anonymous):

70-30=40 \[\frac{30}{40}=\frac{x}{100}\]30*100=40*x 3000=40x x = 3000/40 x=75 percent

OpenStudy (anonymous):

40/30=x/100 4000=30x x=133.33

OpenStudy (ar43r):

30/40*100%=appr. 130%